OpenStudy (anonymous):

write in standard form (0,-4) and (-1,5)

OpenStudy (anonymous):

Yes they are all the same line. You are just saying something like, 2=2 10*2=10*2 (-5+3x+y)+2=(-5+3x+y)+2 Any scalar multiple of your line is the same line. You can always just divide your last equation by -1000 and you will have -3x+y=-8 again. Another nice thing is that if you have 2 equations like: 3x+7y=10 6x+14y=76 If the x and y coefficients are a scalar multiple of each other then the lines are parallel. If the 10 and the 76 were the same number, it would also be the same line. If you take the product of the x coefficients and add them to the product of the 2 y coefficients and it equals 0 the lines are perpendicular. ax+by=# cx+dy=## if (a*c)+(b*d)=0 the lines are perpendicular. Called the dot product for vectors.

OpenStudy (anonymous):

hope this helps:)